# MORPH: Shape-agnostic PDE Foundation Models

### v2

@@ Abstract
We introduce MORPH, a shape-agnostic, autoregressive foundation model for partial differential equations (PDEs). MORPH is built on a convolutional vision transformer backbone that seamlessly handles heterogeneous spatiotemporal datasets of varying data dimensionality (1D–3D) at different resolutions, multiple fields with mixed scalar and vector components. The architecture combines (i) component-wise convolution, which jointly processes scalar and vector channels to capture local interactions, (ii) inter-field cross-attention, which models and selectively propagates information between different physical fields, (iii) axial attentions, which factorizes full spatiotemporal self-attention along individual spatial and temporal axes to reduce computational burden while retaining expressivity. We pretrain multiple model variants on a diverse collection of heterogeneous PDE datasets and evaluate transfer to a range of downstream prediction tasks. Using both full-model fine-tuning and parameter-efficient low-rank adapters (LoRA), MORPH outperforms models trained from scratch in both zero-shot and full-shot generalization. Across extensive evaluations, MORPH matches or surpasses strong baselines and recent state-of-the-art models. Collectively, these capabilities present a flexible and powerful backbone for learning from heterogeneous and multimodal nature of scientific observations, charting a path toward scalable and data-efficient scientific machine learning. The source code, datasets, and models are publicly available at https://github.com/lanl/MORPH.
*Corresponding author: mrautela@lanl.gov
@@ 1 Introduction
Many problems in physics are governed by partial differential equations (PDEs) where field variables exhibit complex spatiotemporal evolution. Machine learning-based surrogate models have made substantial progress in learning such spatiotemporal physical systems. Prominent standalone surrogates include neural operators such as DeepONet (Lu et al., 2021; Kontolati et al., 2024), Fourier Neural Operators (Li et al., 2020), the U-shaped Neural Operator (Rahman et al., 2022) and transformer-based models (Solera-Rico et al., 2024; Li et al., 2022). Inspired by the success of foundation models in natural language processing, recent research directions in scientific machine learning is transitioning from task-specific to task-agnostic modeling. We use the term “PDE foundation model” to denote a universal PDE surrogate or a universal PDE operator pretrained on large, diverse datasets spanning multiple physics. With task-specific fine-tuning, such models can outperform standalone surrogates and demonstrate utility in data- and compute-scarce scenarios.
@@
Partial observability of physics is a bottleneck in learning a PDE foundation model. Two forces collide in practice. Physics: dynamical systems couple nonlinear, multiscale interactions across scalar and vector fields evolving in four-dimensional continuous space-time, often with widely separated characteristic scales. Data: unlike web-scale text or images, pretraining corpora for PDEs are scarce, costly to curate, and frequently provide only partial observability. The existing benchmark spatiotemporal PDE datasets (Takamoto et al., 2022; Ohana et al., 2024; Herde et al., 2024) contain varying data dimensionality (1D–3D) including trajectories, fields, components and resolutions curated through computationally costly simulations. The datasets are generated under restricted parameter settings with limited variation in initial and boundary conditions. In addition, such datasets can be terabytes in size, making naive homogenization or padding prohibitive. On the experimental side, measuring such fields demands substantial investment and dedicated facilities (e.g., wind tunnels for fluid-dynamics experiments). Most measurement modalities also provide partial observability: seismometers, for instance, yield 1D time series on sparse spatial grids, whereas seismic waves propagate as a (3+1)(3{+}1)-D disturbance through Earth’s crust. Similarly, pressure probes and velocimetry typically provide 1D or 2D temporal observations at limited spatial locations, even though the underlying fluid physics exhibits full spatiotemporal fluctuations in the medium. These practical constraints of simulations and measurements motivate a design that can learn effectively from partial information and generalize across data heterogeneity.
@@
A modeling challenge under partial observations. A foundation model with a transformer backbone provides flexibility in learning from partial physical information. Recent PDE foundation models have made substantial progress toward a task-agnostic universal surrogate. However, most of the works (McCabe et al., 2024; Chen et al., 2024; Herde et al., 2024) assume largely homogeneous data dimensionality, with predominantly 2D Cartesian grids, fixed scalar or vector components of field variables, and limited multiphysics coverage. The architectural choices are not designed for data heterogeneity and scale unfavorably as dimensionality and resolution increase often causing memory/compute blow-ups while requiring task-specific reconfiguration during fine-tuning. In particular, their pretraining sets exclude 3D datasets because patching volumetric data expands the token sequence length, driving quadratic self-attention cost $\mathcal{O}(L_{\text{seq}}^{2})$. The same scaling bottleneck discourages inclusion of higher-resolution 2D datasets. Conversely, 1D problems are often excluded because 2D-centric pipelines enforce padding or tiling up to a 2D layout, wasting memory and compute without adding information. Since many experimental measurements are practically 1D (e.g., point probes, line scans, seismometer traces) and provide the most accurate observation of the underlying physics, we regard 1D as a crucial modality. We therefore treat heterogeneity and multimodality as first-class design requirements for a PDE foundation model. Concurrent works like (Ye et al., 2024; McCabe et al., 2024) explicitly highlight the open challenge of supporting more complex domains and extending beyond 2D, underscoring a persistent heterogeneity gap. We identify this as a key bottleneck in current PDE foundation model research. In short, there is a growing need for a flexible and scalable foundation model that can learn from diverse forms of partially observed scientific data.
@@
We introduce MORPH, a PDE foundation model designed to accommodate heterogeneous data across diverse physical phenomenon. MORPH is shape-agnostic, with physics-aware channel handling of PDE datasets. The architecture combines three mechanisms: (a) component-wise convolutions to capture local interactions while jointly processing scalar and vector channels; (b) inter-field cross-attention that models and selectively propagates information across physical fields while condensing them into a single fused representation; and (c) axial attention that factorizes full spatiotemporal self-attention along spatial and temporal axes, reducing compute while retaining expressivity. We define a Unified Physics Tensor Format (UPTF-7) for mini-batches that generalizes across PDE datasets while preserving the semantics of the physical observations. We pretrain MORPH on six heterogeneous spatiotemporal datasets spanning 1D–3D spatial domains with multiple field compositions. We perform full finetuning and parameter-efficient finetuning with low rank adapters (LoRA) on seven additional heterogeneous datasets for downstream prediction tasks. Beyond standard vision transformer hyperparameters, the convolutional and cross-attention modules of the model provide more favorable parameter scaling. We release four model variants—MORPH-Ti (7M), MORPH-S (30M), MORPH-M (126M), and MORPH-L (480M).
@@
  • Shape-agnostic universality. We introduce a shape-agnostic model that operates across arbitrary spatiotemporal PDE dataset dimensionalities (1D–3D), resolutions and mixed scalar/vector fields without task-specific reconfiguration.
  • Expanded context via attention chaos. MORPH employs a larger transformer architecture with one cross-attention and four axial attention modules that attend over a multi-fold increase in spatiotemporal patches (i.e., a larger context window).
@@ 2 Related Work
A wide range of ML-based standalone surrogate models are deployed for solving PDEs. Some of these methods are 3D-CNN (Wandel et al., 2021), ConvLSTM (Shi et al., 2015), DCGAN (Cheng et al., 2020) and GNNs (Chen et al., 2021). Beyond purely data-driven surrogates, physics-informed neural networks (PINNs) has been applied to spatiotemporal problems (Raissi et al., 2019). Recently, neural operators like DeepONet (Goswami et al., 2022), Fourier Neural Operators (FNOs) (Li et al., 2020), wavelet neural operators (WNOs) (Tripura & Chakraborty, 2023), Laplace neural operators (Cao et al., 2024) and physics-informed neural operators (PINOs) (Li et al., 2024b) have gained attention for solving spatiotemporal problems. Another line of methods uses generative models to capture spatiotemporal data distributions. Generative Adversarial Networks (GANs) have been explored to reconstruct or super-resolve turbulent flow fields from partial observations (Buzzicotti et al., 2021). More recently, diffusion models with spatial cross-attention (score-based generative models) have gained traction due to their stability and high fidelity in representing complex distributions (Zhuang et al., 2025; Li et al., 2024a; Gao et al., 2025). Researchers have investigated latent evolution models where spatial learning is performed through autoencoders (Oommen et al., 2022; Vlachas et al., 2022; Maulik et al., 2021; Kontolati et al., 2024; Oommen et al., 2022) and variational autoencoders (Solera-Rico et al., 2024; Rautela et al., 2024). Temporal dynamics in the latent space have been modeled with recurrent neural networks including LSTMs (Maulik et al., 2021; Vlachas et al., 2022; Liu et al., 2022; Rautela et al., 2024), Koopman with non-linear forcing (Khodkar & Hassanzadeh, 2021), DeepONets (Oommen et al., 2022; Kontolati et al., 2024), ODE-Nets (Chen et al., 2018) and transformers with self-attention and easy-attention (Solera-Rico et al., 2024; Alkin et al., 2024).
@@
Standalone PDE surrogates are typically specialized to one equation or parameterized family, which means they must be retrained from scratch whenever the PDE form or conditions change. This lack of generality leads to significant computational overhead without knowledge transfer across PDEs. Some initial works have successfully demonstrated the capability of transfer learning for PDEs (Goswami et al., 2022; Xu et al., 2023). Foundation models for PDEs have emerged to overcome these limitations by capturing the common patterns underlying many different PDE systems. A PDE foundation model is a task-agnostic generalist model pretrained on diverse data followed with minimal task-specific finetuning. Recent works (Yang et al., 2023; Chen et al., 2024; Herde et al., 2024; Chen et al., 2018; Liu et al., 2024; Ye et al., 2024) demonstrated that a PDE foundation model can outperform standalone models. The advantages in generality, data efficiency, transferability, and versatility motivate the development of foundation PDE models over traditional one-off surrogates (Subramanian et al., 2023).
Despite rapid progress, building general-purpose PDE foundation models remains a formidable challenge. PDEs capture complex physics with non-linear interactions, multi-scale spatiotemporal evolutions, discontinuities in the medium (e.g., shocks), and initial and boundary-dependent constraints across multiple fields and physics. The datasets representing such complex spatiotemporal behavior involve several human experts and many compute hours. Public benchmarks remain scarce and frequently provide only partial observability of the underlying physics (Takamoto et al., 2022; Ohana et al., 2024; Herde et al., 2024). It becomes absolutely necessary to build a PDE foundation model with data heterogeneity and modalities in mind (Ye et al., 2024; McCabe et al., 2024). The architectural design should support learning from partial experimental/simulation data and accommodate 1D–3D domains, low to high resolutions, and mixed scalar and vector fields in single and multiphysics settings. However, most of the current PDE foundation models lack this ability.
@@ 3.1 Architecture
Unified Physics Tensor Format (UPTF-7).
The most direct way to handle data heterogeneity is to pad the datasets to a common shape. However, raw PDE datasets can occupy gigabytes to terabytes of storage. For instance, The Well alone contains $\sim$15 TB of physics simulations (Ohana et al., 2024). This makes naive padding both compute and storage-wise inefficient. We therefore seek a unified dataset format that generalizes across PDE datasets while preserving the semantics of the physical observations. We observe that popular benchmarks like PDEBench (Takamoto et al., 2022), The Well (Ohana et al., 2024), pdegym Herde et al. (2024), PDEArena Gupta & Brandstetter (2022), and CFDBench (Yining et al., 2023) can be represented with a generic 7D data shape. We denote it as $(N,T,F,C,D,H,W)$, where $N$ denotes the number of trajectories, $T$ are the number of time steps, $N\times T$ are the number of snapshots, $F$ are the number of field variables, $C$ are the associated scalar/vector components per field, and $(D,H,W)$ are the spatial depth, height, and width. We retain each benchmark’s native on-disk format and convert mini-batches on the fly to our Unified Physics Tensor Format (UPTF-7) when loading to GPUs. For instance, the 2D compressible CFD data in PDEBench contains velocity, pressure, and density fields where velocity is vector-valued while the other two are scalar. Under our convention, scalars are special cases of vectors. Therefore, we broadcast each scalar component to the maximum vector components count when a batch contains mixed scalar–vector fields. Under this convention, the 2D compressible CFD batch has a UPTF-7 layout (Batch,T,F=3,C=2,D=1,H,W)(Batch,T,F{=}3,C{=}2,D{=}1,H,W). Similarly, a 1D Diffusion–Reaction batch with native $(B,T,W)$ layout maps to UPTF-7 as (B,T,F=1,C=1,D=1,H=1,W)(B,T,F{=}1,C{=}1,D{=}1,H{=}1,W), and a 3D MHD batch in The Well maps to (B,T,F=3,C=3,D,H,W)(B,T,F{=}3,C{=}3,D,H,W). We defer more dataset-specific details to Appendix B.2.1.
Overview of MORPH.
@@
The convolutional operator acts on the component channel(s), producing a set of feature maps determined by the number of filters. Adding convolutional inductive bias helps vision transformers capture locality and improves sample efficiency (Dosovitskiy et al., 2020). We can scale feature maps and ultimately MORPH via the number of convolutional filters (default: 8). We partition feature maps into patches of size 8. Across pretraining and fine-tuning datasets, we use at most 4,096 patches, yielding a 32,768-dimensional feature vector, which we linearly project to an embedding whose dimensionality ranges from 256 to 1,024 depending on the model variant.
@@
In most PDE datasets, multiple fields with different spatial scales evolve at various temporal rates. Field-wise multi-head cross-attention prioritizes relevant inter-field interactions and selectively propagates information, enabling careful feature fusion while suppressing spurious activations arising from scale mismatches among coupled fields. The cross-attention operator outputs a single fused field, significantly reducing computational cost (Nguyen et al., 2023). We scale MORPH further via the cross-attention dimension, which by default matches those of axial self-attention.
@@
We factorize full 4D spatiotemporal attention for a (3+1)D dynamical system into time-wise, depth-wise, height-wise, and width-wise axial attentions. For high-dimensional datasets, this limits the computational cost to $O(T^{2}+D^{2}+H^{2}+W^{2})$, in contrast to full spatiotemporal attention with cost $O((TDHW)^{2})$ (Ho et al., 2019). Time-wise axial attention supports flexible temporal context allowing fixed or variable order autoregression (default: 1 step). More details on modeling aspects are deferred to Appendix B.1.
3.2 Data preprocessing
Normalization.
We use Reversible Instance Normalization (ReVIN) to counter covariate shift across heterogeneous datasets and training regimes. It was first proposed for time-series forecasting and validated to mitigate distribution shift (Kim et al., 2021). We normalize the data and cache the corresponding means and standard deviations, then use these statistics to exactly denormalize model outputs. Unlike McCabe et al. (2024), which normalizes on-the-fly, we pre-normalize the entire dataset and train on normalized batches, reducing the normalization overhead incurred during training and fine-tuning.
Data streaming and sharding setup.
MORPH is designed to handle data heterogeneity and diversity, which introduces system-level challenges. To utilize CPU, GPU, and RAM efficiently, we make several non-traditional choices. One issue is that computing cluster resources impose RAM limits, making it difficult to load all datasets into memory simultaneously. As PDE benchmarks continue to grow, streaming becomes essential for building large PDE foundation models. Accordingly, we stream datasets from storage in chunks, which uses RAM efficiently and scales with the number and size of datasets. Another issue is the wide range of token lengths (128–4096) across pretraining datasets, which leads to varying computational burdens in parallel processing across batches. Instead of a single PyTorch DataLoader for all datasets, we use one dataloader per dataset. With six pretraining datasets, we run six DataLoaders. This enables dataset-specific choices of batch sizes and number of workers. Given this multi-dataloader setup, data sharding under DistributedDataParallel(DDP) is nontrivial, so we implement custom sharding across workers and ranks. We defer more details to Appendix B.2.2.

### v3

# MORPH: PDE Foundation Models with Arbitrary Data Modality

@@ Abstract
We introduce MORPH, a modality-agnostic, autoregressive foundation model for partial differential equations (PDEs). MORPH is built on a convolutional vision transformer backbone that seamlessly handles heterogeneous spatiotemporal datasets of varying data modality (1D–3D) at different resolutions, and multiple fields with mixed scalar and vector components. The architecture combines (i) component-wise convolution, which jointly processes scalar and vector channels to capture local interactions, (ii) inter-field cross-attention, which models and selectively propagates information between different physical fields, (iii) axial attentions, which factorize full spatiotemporal self-attention along individual spatial and temporal axes to reduce computational burden while retaining expressivity. We pretrain multiple model variants on a diverse collection of heterogeneous PDE datasets and evaluate transfer to a range of downstream prediction tasks. Using both full-model fine-tuning and parameter-efficient low-rank adapters (LoRA), MORPH outperforms models trained from scratch. Across extensive evaluations, MORPH matches or surpasses strong baselines and recent state-of-the-art models. Collectively, these capabilities present a flexible and powerful backbone for learning from the heterogeneous and multimodal nature of scientific observations, charting a path toward scalable and data-efficient scientific machine learning. The source code, datasets, and models are publicly available at https://github.com/lanl/MORPH.
@@ 1 Introduction
Many problems in physics are governed by partial differential equations (PDEs) where field variables exhibit complex spatiotemporal evolution. Machine learning-based surrogate models have made substantial progress in learning such spatiotemporal physical systems. Prominent standalone surrogates include neural operators such as DeepONet (lu2021learning; kontolati2024learning), Fourier Neural Operators (li2020fourier), the U-shaped Neural Operator (rahman2022u) and transformer-based models (solera2024beta; li2022transformer). Inspired by the success of foundation models in natural language processing, recent research directions in scientific machine learning are transitioning from task-specific to task-agnostic modeling. We use the term “PDE foundation model” to denote a universal PDE surrogate or a universal PDE operator pretrained on large, diverse datasets spanning multiple physics. With task-specific fine-tuning, such models can outperform standalone surrogates and demonstrate utility in data- and compute-scarce scenarios.
@@
Partial observability of physics is a bottleneck in learning a PDE foundation model. Two forces collide in practice. Physics: dynamical systems couple nonlinear, multiscale interactions across scalar and vector fields evolving in four-dimensional continuous space-time, often with widely separated characteristic scales. Data: unlike web-scale text or images, pretraining corpora for PDEs are scarce, costly to curate, and frequently provide only partial observability. The existing benchmark spatiotemporal PDE datasets (takamoto2022pdebench; ohana2024well; herde2024poseidon) contain varying data modality (1D–3D) including trajectories, fields, components and resolutions curated through computationally costly simulations. The datasets are generated under restricted parameter settings with limited variation in initial and boundary conditions. In addition, such datasets can be terabytes in size, making naive homogenization or padding prohibitive. On the experimental side, measuring such fields demands substantial investment and dedicated facilities (e.g., wind tunnels for fluid-dynamics experiments). Most measurement modalities also provide partial observability: seismometers, for instance, yield 1D time series on sparse spatial grids, whereas seismic waves propagate as a (3+1)(3{+}1)-D disturbance through Earth’s crust. Similarly, pressure probes and velocimetry typically provide 1D or 2D temporal observations at limited spatial locations, even though the underlying fluid physics exhibits full spatiotemporal fluctuations in the medium. These practical constraints of simulations and measurements motivate a design that can learn effectively from partial information and generalize across data heterogeneity.
@@
A modeling challenge under partial observations. A foundation model with a transformer backbone provides flexibility in learning from partial physical information. Recent PDE foundation models have made substantial progress toward a task-agnostic universal surrogate. However, most of the works (mccabe2024multiple; chen2024data; herde2024poseidon) assume largely homogeneous data modality, with predominantly 2D Cartesian grids, fixed scalar or vector components of field variables, and limited multiphysics coverage. The architectural choices are not designed for data heterogeneity and scale unfavorably as data modality changes and resolution increases often causing memory/compute blow-ups while requiring task-specific reconfiguration during fine-tuning. In particular, their pretraining sets exclude 3D datasets because patching volumetric data expands the token sequence length, driving quadratic self-attention cost $\mathcal{O}(L_{\text{seq}}^{2})$. The same scaling bottleneck discourages inclusion of higher-resolution 2D datasets. Conversely, 1D problems are often excluded because 2D-centric pipelines enforce padding or tiling up to a 2D layout, wasting memory and compute without adding information. Since many experimental measurements are practically 1D (e.g., point probes, line scans, seismometer traces) and provide the most accurate observation of the underlying physics, we regard 1D as a crucial modality. We therefore treat heterogeneity and multimodality as first-class design requirements for a PDE foundation model. Concurrent works like (ye2024pdeformer; mccabe2024multiple) explicitly highlight the open challenge of supporting more complex domains and extending beyond 2D, underscoring a persistent heterogeneity gap. We identify this as a key bottleneck in current PDE foundation model research. In short, there is a growing need for a flexible and scalable foundation model that can learn from diverse forms of partially observed scientific data.
@@
We introduce MORPH, a PDE foundation model designed to accommodate heterogeneous data across diverse physical phenomena. MORPH is shape-agnostic, with physics-aware channel handling of PDE datasets. The architecture combines three mechanisms: (a) component-wise convolutions to capture local interactions while jointly processing scalar and vector channels; (b) inter-field cross-attention that models and selectively propagates information across physical fields while condensing them into a single fused representation; and (c) axial attention that factorizes full spatiotemporal self-attention along spatial and temporal axes, reducing compute while retaining expressivity. We define a Unified Physics Tensor Format (UPTF-7) for mini-batches that generalizes across PDE datasets while preserving the semantics of the physical observations. We pretrain MORPH on six heterogeneous spatiotemporal datasets spanning 1D–3D spatial domains with multiple field compositions. We perform full finetuning and parameter-efficient finetuning with low rank adapters (LoRA) on seven additional heterogeneous datasets for downstream prediction tasks. We release four model variants: MORPH-Ti (7M), MORPH-S (30M), MORPH-M (126M), and MORPH-L (480M).
@@
  • Modality-agnostic universality. We introduce a modality-agnostic model that operates across arbitrary spatiotemporal PDE dataset dimensionalities (1D–3D), resolutions and mixed scalar/vector fields without task-specific reconfiguration.
  • Expanded spatiotemporal context. MORPH employs a larger transformer architecture with one cross-attention and four axial attention modules that attend over a multi-fold increase in spatiotemporal patches (i.e., a larger context window).
@@ 2 Related Work
A wide range of ML-based standalone surrogate models are deployed for solving PDEs. Some of these methods are 3D-CNN (wandel2021teaching), ConvLSTM (shi2015convolutional), DCGAN (cheng2020data) and GNNs (chen2021graph). Beyond purely data-driven surrogates, physics-informed neural networks (PINNs) have been applied to spatiotemporal problems (RAISSI2019686). Recently, neural operators like DeepONet (goswami2022deep), Fourier Neural Operators (FNOs) (li2020fourier), wavelet neural operators (WNOs) (tripura2023wavelet), Laplace neural operators (cao2024laplace) and physics-informed neural operators (PINOs) (li2024physics) have gained attention for solving spatiotemporal problems. Another line of methods uses generative models to capture spatiotemporal data distributions. Generative Adversarial Networks (GANs) have been explored to reconstruct or super-resolve turbulent flow fields from partial observations (buzzicotti2021reconstruction). More recently, diffusion models with spatial cross-attention (score-based generative models) have gained traction due to their stability and high fidelity in representing complex distributions (zhuang2025spatially; li2024learning; gao2025generative). Researchers have investigated latent evolution models where spatial learning is performed through autoencoders (oommen2022learning; vlachas2022multiscale; maulik2021reduced; kontolati2024learning) and variational autoencoders (solera2024beta; rautela2024conditional). Temporal dynamics in the latent space have been modeled with recurrent neural networks including LSTMs (maulik2021reduced; vlachas2022multiscale; liu2022hierarchical; rautela2024conditional), Koopman with non-linear forcing (khodkar2021data), DeepONets (oommen2022learning; kontolati2024learning), ODE-Nets (chen2018neural) and transformers with self-attention and easy-attention (solera2024beta; alkin2024universal).
@@
Standalone PDE surrogates are typically specialized to one equation or parameterized family, which means they must be retrained from scratch whenever the PDE form or conditions change. This lack of generality leads to significant computational overhead without knowledge transfer across PDEs. Some initial works have successfully demonstrated the capability of transfer learning for PDEs (goswami2022deep; xu2023transfer). Foundation models for PDEs have emerged to overcome these limitations by capturing the common patterns underlying many different PDE systems. A PDE foundation model is a task-agnostic generalist model pretrained on diverse data, followed with minimal task-specific finetuning. Recent works (yang2023context; chen2024data; herde2024poseidon; chen2018neural; liu2024prose; ye2024pdeformer) demonstrated that a PDE foundation model can outperform standalone models. The advantages in generality, data efficiency, transferability, and versatility motivate the development of foundation PDE models over traditional one-off surrogates (subramanian2023towards). Despite rapid progress, building general-purpose PDE foundation models remains a formidable challenge. PDEs capture complex physics and the datasets representing such spatiotemporal behavior involve several human experts and many compute hours. Public benchmarks remain scarce and frequently provide only partial observability of the underlying physics (takamoto2022pdebench; ohana2024well; herde2024poseidon). It becomes absolutely necessary to build a PDE foundation model with data heterogeneity and modalities in mind (ye2024pdeformer; mccabe2024multiple; morel2025disco). The architectural design should support learning from partial experimental/simulation data and accommodate 1D–3D domains, low to high resolutions, and mixed scalar and vector fields in single and multiphysics settings. However, most of the current PDE foundation models lack this ability.
@@ 3 Proposed Method
3.1 Unified Physics Tensor Format (UPTF-7)
The most direct way to handle data heterogeneity is to pad the datasets to a common shape. However, raw PDE datasets can occupy gigabytes to terabytes of storage. For instance, The Well alone contains $\sim$15 TB of physics simulations (ohana2024well). This makes naive padding both compute and storage-wise inefficient. We therefore seek a unified dataset format that generalizes across PDE datasets while preserving the semantics of the physical observations. We observe that popular benchmarks like PDEBench (takamoto2022pdebench), The Well (ohana2024well), PDEGym (herde2024poseidon), PDEArena (gupta2022towards), and CFDBench (CFDBench) can be represented with a generic 7D data shape. We denote it as $(N,T,F,C,D,H,W)$, where $N$ denotes the number of trajectories, $T$ are the number of time steps, $N\times T$ are the number of snapshots, $F$ are the number of field variables, $C$ are the associated scalar/vector components per field, and $(D,H,W)$ are the spatial depth, height, and width. We retain each benchmark’s native on-disk format and convert mini-batches on the fly to our Unified Physics Tensor Format (UPTF-7) when loading to GPUs. For instance, the 2D compressible CFD data in PDEBench contains velocity, pressure, and density fields where velocity is vector-valued while the other two are scalar. Under our convention, scalars are special cases of vectors. Therefore, we broadcast each scalar component to the maximum vector components count when a batch contains mixed scalar–vector fields. Therefore, the 2D compressible CFD batch maps to a UPTF-7 layout (B,T,F=3,C=2,D=1,H,W)(B,T,F{=}3,C{=}2,D{=}1,H,W). Similarly, a 1D Diffusion–Reaction batch with native $(B,T,W)$ layout maps to UPTF-7 as (B,T,F=1,C=1,D=1,H=1,W)(B,T,F{=}1,C{=}1,D{=}1,H{=}1,W), and a 3D MHD batch in The Well maps to (B,T,F=3,C=3,D,H,W)(B,T,F{=}3,C{=}3,D,H,W). We defer more dataset-specific details to Appendix A.4.
3.2 Overview of MORPH.
@@ 3.1 Unified Physics Tensor Format (UPTF-7)
The convolutional operator acts on the component channel(s), producing a set of feature maps determined by the number of filters. Adding convolutional inductive bias helps vision transformers capture locality and improves sample efficiency (dosovitskiy2020image). We can scale feature maps and ultimately MORPH via the number of convolutional filters (default: 8). We partition feature maps into patches of size 8. Across pretraining and fine-tuning datasets, we use at most 4,096 patches, yielding a 32,768-dimensional feature vector, which we linearly project to an embedding whose dimensionality ranges from 256 to 1,024 depending on the model variant.
@@
In most PDE datasets, multiple fields with different spatial scales evolve at various temporal rates. Field-wise multi-head cross-attention prioritizes relevant inter-field interactions and selectively propagates information, enabling careful feature fusion while suppressing spurious activations arising from scale mismatches among coupled fields. The cross-attention operator outputs a single fused field, significantly reducing computational cost (nguyen2023climax). We scale MORPH further via the cross-attention dimension, which by default matches those of axial self-attention.
@@
We factorize full 4D spatiotemporal attention for a (3+1)D dynamical system into time-wise, depth-wise, height-wise, and width-wise axial attentions. For high-dimensional datasets, this limits the computational cost to $O(T^{2}+D^{2}+H^{2}+W^{2})$, in contrast to full spatiotemporal attention with cost $O((TDHW)^{2})$ (ho2019axial). Time-wise axial attention supports flexible temporal context allowing fixed or variable order autoregression (default: 1 step). More details on modeling aspects are deferred to Appendix B.1.
3.3 Transfer across data modalities
A central question for a PDE foundation model is whether representations learned on one modality (e.g., 2D incompressible flows) transfer across multiple data modalities such as 1D, 2D and 3D. MORPH is explicitly designed to be modality-agnostic while maintaining generalization across diverse physics. To test this hypothesis, we pretrain MORPH exclusively on the 2D incompressible Navier–Stokes dataset (CFD2D-IC) under an autoregressive objective. We then perform zero-shot inference (no fine-tuning, all weights frozen) on six heterogeneous targets (fluids based) spanning different modalities: CFD1D, CFD2D, CFD3D, CFD3D-Turb, MHD3D, and TGC3D.
To quantify transfer, we define zero-shot Gap-Closure Ratio (GCR), which measures what fraction of the performance gap between a naive, training-free baseline and a fully trained-from-scratch model is closed by the pretrained model in the zero-shot setting. Formally,
[…]
Here, $E_{\mathrm{PT}}$ is the zero-shot error of the CFD2D-IC pretrained model evaluated on a given target, $E_{\mathrm{SC}}$ is the error of an identical architecture trained fully from scratch on that target, and $E_{\mathrm{NB}}$ is the error of a training-free naive baseline (the same architecture with random initialization and no learning, informally corresponding to near–coin-toss performance), taken as the minimum over 25 random seeds. We define $1-E_{\mathrm{PT}}/E_{\mathrm{NB}}$ as the naive-baseline gain of the pretrained model (NBG-PT), and $1-E_{\mathrm{SC}}/E_{\mathrm{NB}}$ as the naive-baseline gain of the trained-from-scratch model (NBG-SC). In simpler terms, GCR is the ratio $\mathrm{NBG\text{-}PT}$ and $\mathrm{NBG\text{-}SC}$. Our definition of zero-shot GCR is inspired by performance-gap-recovered (PGR) metric introduced in burns2023weak.
Interpretation: $\mathrm{GCR}=0$ means zero-shot does no better than naive. $\mathrm{GCR}=1$ means zero-shot matches a fully trained scratch model; $\mathrm{GCR}>1$ indicates zero-shot surpasses a model trained from scratch; any $\mathrm{GCR}>0$ demonstrates positive transfer across modalities and physics. Fig. 2 summarizes transfer from CFD2D-IC to the six targets via three quantities: NBG-PT, NBG-SC and the resulting $\mathrm{GCR}$. We observe consistent positive transfer ($\mathrm{GCR}>0$) across all fluid-specific targets (green bar), with the strongest effects on near-domain problems (e.g., compressible CFD1D, CFD2D and CFD3D) and attenuated but still nontrivial transfer to CFD3D (with turbulence) and multiphysics settings such as MHD3D and self-gravitating TGC3D. These results indicate that 2D pretraining learns reusable operators that carry to 1D/2D/3D data modalities, providing empirical support for MORPH’s modality-agnostic design.
Figure 2: Zero-shot transfer from CFD2D-IC pretraining. Targets: CFD1D, CFD2D, CFD3D, MHD3D, CFD3D-Turb, TGC3D. Bars show the Naive Baseline Gain of the pretrained model (NBG-PT), the Naive Baseline Gain of a trained-from-scratch model (NBG-SC), and the resulting Gap-Closure Ratio (GCR). $\mathrm{GCR}>0$ indicates cross-modality transfer.